The sides of a triangle are 7cm, 8cm and 5cm respectively. Find its area
Found 2 solutions by ikleyn, rothauserc:
Answer by ikleyn(52847)

.
Use the Heron's formula.

---------------------
On Heron's formula see the lessons
    - Proof of the Heron's formula for the area of a triangle
    - One more proof of the Heron's formula for the area of a triangle
    - Solved problems on area of triangles
in this site.

Also,  you have this free of charge online textbook on Geometry
    GEOMETRY - YOUR ONLINE TEXTBOOK
in this site.

The referred lessons are the part of this online textbook under the topic "Area of triangles".

Save the link to this online textbook together with its description

Free of charge online textbook in GEOMETRY
https://www.algebra.com/algebra/homework/Triangles/GEOMETRY-your-online-textbook.lesson

to your archive and use it when it is needed.



Answer by rothauserc(4718)

Use Heron's formula
:
Area (A) = square root(s(s-a)(s-b)(s-c)) where s = (a + b + c)/2 and a, b, c are the lengths of the sides of the triangle
:
s = (7 + 8 + 5)/2 = 10
:
A = square root(10(10-7)(10-8)(10-5)) = square root(300) = 10square root(3) cm^2
:
Note 10square root(3) is approximately 17.32 cm^2, where cm^2 is square centimeters
: